# Gravitational effects on fluid dynamics in cosmological first-order phase transitions

### v2

@@ 3 Fluid dynamics in FOPTs beyond the self-similarity
In this section, we introduce the cosmological hydrodynamic scheme proposed in Ref. [82], which solves the peculiar motion of fluid in a (nearly) Friedmann-Lemaître-Robertson-Walker (FLRW) spacetime. Then we explain how to incorporate the Higssless scheme into this scheme to investigate the fluid dynamics around the slowly expanding bubble wall.
@@ 3.1 Hydrodynamic equations around the FLRW spacetime
We assume the following spacetime metric following Ref. [82], in which the peculiar motion of relativistic fluid around the FLRW spacetime is considered
$ds^{2}$ $=-(1+2\Phi(t,r))dt^{2}-2\chi(t,r)dtdr+(1-2\Psi(t,r))a(t)^{2}(dr^{2}+r^{2}d\Omega^{2}).$ (3.1)
Since we are interested in the fluid dynamics around a single bubble, spherical symmetry is assumed here. Note that, while $\chi$ does not play an essential role in the following discussion, it must be included in deriving the equations of motion [85]. After taking the uniform-expansion gauge that assumes vanishing perturbed part of the trace of extrinsic curvature, the equations of motion for the fluid and metric perturbations are derived in Ref. [82]. With the fluid four-velocity $u^{\mu}=\gamma(1,v(t,r)/a,0,0)$, we find the following equations of motion:
@@ 3.2 Application to FOPTs
In this case, the background scale factor $a(t)$ shows relatively complicated behavior. The Friedmann equation
@@
where $H_{0}$ is the initial Hubble parameter at $t=0$ and $\alpha_{0}=\epsilon/\rho_{\rm rad}(t=0)$. Throughout this paper, we assume that the true-vacuum bubble nucleates during the radiation domination ($\alpha_{0}\ll 1$), which is actually important for satisfying the weak backreaction condition. In App. A, we summarize the alternative formulation and the results for the case where only radiation is taken as background. Though there exists a certain complication in this decomposition compared to the above one, we found that the different background settings do not change our main results.
@@
$\Delta\Phi(t)\equiv|\Phi(t,0)-\Phi(t,r\to\infty)|\simeq\frac{3}{8M_{\rm Pl}^{2}}w_{b}(0)\alpha_{0}a^{2}(t)r^{2}_{w}(t).$
@@
One also needs to care about the short-scale condition. The last condition in Eq. (3.5) is derived from the perturbativity of $\chi(t,r)$ [82]
$\frac{\partial_{r}\chi}{a^{2}}$ $\ll v,$ (3.23)
with the ADM momentum constraint equation $\partial_{r}^{3}\chi/a^{4}\sim w\gamma^{2}v/M_{\rm pl}^{2}$. This implies that the physical length scale $l_{c}$ should be estimated from the region where the fluid velocity $v$ is nonzero. As long as the perturbativity condition for $\Phi$ is satisfied, it is safely taken to be the order of the shell thickness in the flat background, which is much smaller than the bubble size and the Hubble horizon. However, once $\Phi$ becomes of ${\cal O}(0.1)$, fluid velocity starts to be induced over the Hubble size, and the system size can no longer be regarded as the shell thickness. Since this transition is triggered by $\Phi\sim{\cal O}(0.1)$, we can monitor the violation of the short scale condition by keeping track of the perturbativity of $\Phi$.
@@ 4.1.2 Late time solutions and the efficiency factor
The results of the simulation are shown in Fig 4. In this figure, we solve the system for the same parameter values as in Fig 2. To evaluate $\Delta\Phi(t)$, we increase the box size by lowering the spatial resolution to $[\Delta r]_{\rm sim}=10^{-3}H_{0}^{-1}$ while keeping $n_{r}=20000$. We also set $[\Delta t]_{\rm sim}=5\times 10^{-4}H_{0}^{-1}$ to facilitate a longer time scale in the simulation.
Note that the late-time behavior of the system is not much affected by the reduced resolution. This is because the late-time behavior is dominantly determined by the late-time energy injection: though the spatial resolution at initial times is lower than what is needed to fully capture the shell profile, the error at these times is negligible. The asymptotic profile is determined by the late-time energy injection by large $\alpha(t)$, when the bubble and fluid shell has grown large enough to be captured by the current spatial resolution. Here we also add the luminal case $v_{w}=1$ since the wall velocity typically reaches to the speed of light in the scenarios of extremely supercooled FOPT.
The late-time results are shown in Fig 4. We observe an interesting time evolution in the third example where $v_{w}=0.8$. The fluid profile is in the detonation regime according to $(\alpha_{0},v_{w})=(10^{-2},0.8)$ at earlier times, while at later times the larger value of $\alpha(t)$ drives the fluid profile into the hybrid regime. This behavior is indeed expected from the phase diagram of flat spacetime solutions depicted in, for example, Fig. 7 of Ref. [72]. When $v_{w}>c_{s}$, as $\alpha$ is increased from a small value to a large value, one can see that the profile transitions from the detonation region to the hybrid region. As $\alpha(t)$ increases with time in our case, such a transition is expected to inevitably occur depending on the initial conditions. Therefore, we conclude that such a transition from detonation to hybrid solution is a general behavior for the cosmological time scale evolution.
@@
We also found that the late time magnitude of $\kappa(t)$ (and also $\Delta\Phi(t)$) seems to converge when $\alpha_{0}$ decreases, by comparing Figs. 6 and 7. To investigate this behavior, we also simulated the system for $\alpha_{0}=10^{-4}$ with specific values of $v_{w}=0.3,0.5,0.7,0.9$. In Fig. 8, the evolution of $\kappa(t)$ is shown for different wall velocities and $\alpha_{0}=10^{-1},10^{-2},10^{-3},10^{-4}$. One can understand that the larger value of $\kappa(t)$ for larger $\alpha_{0}$ is because the shell is still in the process of becoming thinner. In either wall velocity, the degree of decrease in the value of $\kappa(t)$ becomes smaller as $\alpha_{0}$ decreases. This could be consistent with Ref. [76] that derives the late-time limit solution, indicating the existence of a limit to the thinning of fluid shell. We believe that this convergent behavior is reasonable since for $\alpha_{0}$ small enough, late time dynamics driven by the horizon-scale bubble should become independent of precisely when the bubble nucleates. Here again we plot the correspondent value of $\kappa$ for the flat spacetime solutions and one can read to what extent $\kappa(t)$ is suppressed over the cosmological evolution. [^6] For $v_{w}=0.5$ and $v_{w}=0.7$, the shock profile is not completely captured in the earliest time steps and this results in a slight deviation from the flat spacetime case. Notice that this is not a physical deviation, but the deviation in the later time steps is certainly caused by the cosmological expansion. For interested readers, we also present the similar plot for different physical quantities in App. B.
@@ Acknowledgments
The authors thank Naoki Yoshida, Thomas Konstandin, Isak Stomberg, Lorenzo Giombi, Mark Hindmarsh, and Henrique Rubira for fruitful discussions. The work of RJ is supported by JSPS KAKENHI Grant Numbers 23K17687 and 23K19048. JK is supported by the JSPS Overseas Research Fellowships.
Appendix A True vacuum region as the background
@@
$\delta\tilde{p}$ $=\frac{1}{4}\delta\tilde{w}-\frac{3}{4}\alpha(t)\Theta(r-r_{w}(t)),$ (A.2)
$\delta\tilde{\rho}$ $=\frac{3}{4}\delta\tilde{w}+\frac{3}{4}\alpha(t)\Theta(r-r_{w}(t)).$ (A.3)
@@
Appendix B Late time evolution

### v3

@@ 3 Fluid dynamics in FOPTs beyond the self-similarity
In this section, we introduce the cosmological hydrodynamic scheme proposed in Ref. [82], which solves the peculiar motion of fluid with relativistic pressure in a (nearly) Friedmann-Lemaître-Robertson-Walker (FLRW) spacetime. Then we explain how to incorporate the Higssless scheme into this scheme to investigate the fluid dynamics around the slowly expanding bubble wall.
@@ 3.1 Hydrodynamic equations around the FLRW spacetime
We assume the following spacetime metric following Ref. [82], in which the peculiar motion of relativistic fluid around the FLRW spacetime is considered:
$ds^{2}$ $=-(1+2\Phi(t,r))dt^{2}-2\partial_{r}\chi(t,r)dtdr+(1-2\Psi(t,r))a(t)^{2}(dr^{2}+r^{2}d\Omega^{2}).$ (3.1)
Since we are interested in the fluid dynamics around a single bubble, spherical symmetry is assumed here. After taking the uniform-expansion gauge (UEG) that assumes vanishing perturbed part of the trace of extrinsic curvature, the equations of motion for the fluid with relativistic pressure and metric perturbations are derived in Ref. [82].
With the fluid four-velocity $u^{\mu}=\gamma(1,v(t,r)/a,0,0)$, we find the following equations of motion:
@@
Before proceeding, we comment on the gauge choice in the present paper. As mentioned in the beginning of this section, we follow Ref. [82] and adopt the UEG, in which the trace of the extrinsic curvature $\kappa$ is set to zero. This choice, together with the spatial gauge condition, fixes the gauge completely and leaves no further gauge degrees of freedom. Consequently, all the remaining variables can be regarded as gauge invariant quantities, ensuring the absence of gauge artifacts [85, 86]. Notice that the UEG is not a common choice in the context of cosmological perturbation theory. For interested readers, we provide a brief discussion in App. A, where we compare physical quantities in the UEG and in the zero-shear gauge (ZSG), which is more standard in this context. In the regime of interest, the differences between the quantities in UEG and ZSG are expected to be suppressed by a common factor related to the thickness of fluid shell.
Despite the absence of gauge artifacts, one may question whether the quantities presented in this paper (e.g., fluid velocity and enthalpy) are suitable for direct comparison with their counterparts in flat spacetime. That is, it might be possible that quantities in one gauge is more suited than another for comparison with quantities in the flat background. This concern reflects a broader issue in general relativity, where the meaning of physical quantities is inherently tied to the gauge choice. In this regard, we note that the UEG is the one that makes the expansion rate uniform over each spatial slice. This property isolates local fluid motions from spatial variations in the expansion rate that would be present in other gauges, providing a clearer interpretation of fluid dynamics in an expanding spacetime. Thus, we conclude that the present gauge choice is at least well-suited, albeit not proven to be the best, for comparisons with results in flat spacetime.
@@ 3.2 Application to FOPTs
In this case, the background scale factor $a(t)$ shows a relatively complicated behavior. The Friedmann equation
@@
where $H_{0}$ is the initial Hubble parameter at $t=0$ and $\alpha_{0}=\epsilon/\rho_{\rm rad}(t=0)$. Throughout this paper, we assume that the true-vacuum bubble nucleates during the radiation domination ($\alpha_{0}\ll 1$), which is actually important for satisfying the weak backreaction condition. In App. B, we summarize the alternative formulation and the results for the case where only radiation is taken as background. Though there exists a certain complication in this decomposition compared to the above one, we found that the different background settings do not change our main results.
@@
$\Delta\Phi(t)$ $\equiv|\Phi(t,0)-\Phi(t,r\to\infty)|\simeq\frac{3}{8M_{\rm Pl}^{2}}w_{b}(0)\alpha_{0}a^{2}(t)r^{2}_{w}(t).$ (3.23)
@@
One also needs to care about the short-scale condition. The last condition in Eq. (3.5) is actually connected to the perturbativity of $\chi(t,r)$ [82]
$\frac{\partial_{r}\chi}{a}$ $\ll v,$ (3.24)
which is derived with the ADM momentum constraint equation $\partial_{r}^{3}\chi/a^{3}\sim w\gamma^{2}v/M_{\rm pl}^{2}$. This implies that the physical length scale $l_{c}$ should be estimated from the region where the fluid velocity $v$ is nonzero. As long as the perturbativity condition for $\Phi$ is satisfied, it is safely taken to be the order of the shell thickness in the flat background, which is much smaller than the bubble size and the Hubble horizon. However, once $\Phi$ becomes of ${\cal O}(0.1)$, fluid velocity starts to be induced over the Hubble size, and the system size can no longer be regarded as the shell thickness. Since this transition is triggered by $\Phi\sim{\cal O}(0.1)$, we can monitor the violation of the short scale condition by keeping track of the perturbativity of $\Phi$.
@@ 4.1.2 Late time solutions and the efficiency factor
The results of the simulation are shown in Fig 4. In this figure, we solve the system for the same parameter values as in Fig 2. To evaluate $\Delta\Phi(t)$, we increase the box size by lowering the spatial resolution to $[\Delta r]_{\rm sim}=10^{-3}H_{0}^{-1}$ while keeping $n_{r}=20000$. We also set $[\Delta t]_{\rm sim}=5\times 10^{-4}H_{0}^{-1}$ to facilitate a longer time scale in the simulation. Note that the late-time behavior of the system is not much affected by the reduced resolution. This is because the late-time behavior is dominantly determined by the late-time energy injection: though the spatial resolution at initial times is lower than what is needed to fully capture the shell profile, the error at these times is negligible. The asymptotic profile is determined by the late-time energy injection by large $\alpha(t)$, when the bubble and fluid shell has grown large enough to be captured by the current spatial resolution. Here we also add the luminal case $v_{w}=1$ since the wall velocity typically reaches to the speed of light in the scenarios of extremely supercooled FOPT.
We observe an interesting time evolution in the third example of Fig 4 where $v_{w}=0.8$. The fluid profile is in the detonation regime according to $(\alpha_{0},v_{w})=(10^{-2},0.8)$ at earlier times, while at later times the larger value of $\alpha(t)$ drives the fluid profile into the hybrid regime. This behavior is indeed expected from the phase diagram of flat spacetime solutions depicted in, for example, Fig. 7 of Ref. [72]. When $v_{w}>c_{s}$, as $\alpha$ is increased from a small value to a large value, one can see that the profile transitions from the detonation region to the hybrid region. As $\alpha(t)$ increases with time in our case, such a transition is expected to inevitably occur depending on the initial conditions. Therefore, we conclude that such a transition from detonation to hybrid solution is a general behavior for the cosmological time scale evolution.
@@
We also found that the late time magnitude of $\kappa(t)$ (and also $\Delta\Phi(t)$) seems to converge when $\alpha_{0}$ decreases, by comparing Figs. 6 and 7. To investigate this behavior, we also simulated the system for $\alpha_{0}=10^{-4}$ with specific values of $v_{w}=0.3,0.5,0.7,0.9$. In Fig. 8, the evolution of $\kappa(t)$ is shown for different wall velocities and $\alpha_{0}=10^{-1},10^{-2},10^{-3},10^{-4}$. One can understand that the larger value of $\kappa(t)$ for larger $\alpha_{0}$ is because the shell is still in the process of becoming thinner. In either wall velocity, the degree of decrease in the value of $\kappa(t)$ becomes smaller as $\alpha_{0}$ decreases. This could be consistent with Ref. [76] that derives the late-time limit solution, indicating the existence of a limit to the thinning of fluid shell. We believe that this convergent behavior is reasonable since for $\alpha_{0}$ small enough, late time dynamics driven by the horizon-scale bubble should become independent of precisely when the bubble nucleates. Here again we plot the correspondent value of $\kappa$ for the flat spacetime solutions and one can read to what extent $\kappa(t)$ is suppressed over the cosmological evolution. [^6] For $v_{w}=0.5$ and $v_{w}=0.7$, the shock profile is not completely captured in the earliest time steps and this results in a slight deviation from the flat spacetime case. Notice that this is not a physical deviation, but the deviation in the later time steps is certainly caused by the cosmological expansion. For interested readers, we also present the similar plot for different physical quantities in App. C.
@@ Acknowledgments
The authors thank Naoki Yoshida, Thomas Konstandin, Isak Stomberg, Lorenzo Giombi, Mark Hindmarsh, and Henrique Rubira for fruitful discussions. The work of RJ is supported by JSPS KAKENHI Grant Numbers 23K17687, 23K19048, and 24K07013. JK is supported by the JSPS Overseas Research Fellowships. JK acknowledges support from Istituto Nazionale di Fisica Nucleare (INFN) through the Theoretical Astroparticle Physics (TAsP) project.
Appendix A Comparison with a commonly used gauge choice
In the cosmological perturbation theory, one common choice is the zero-shear gauge (ZSG), in which the temporal gauge condition is chosen such that $\chi=0$ instead of $\kappa=0$. In this gauge, the variables $\Phi$ and $\Psi$ correspond to the well-known Bardeen potentials [92]. While the gauge-invariant scalar potential of velocity perturbation $v_{0}(t,r)$ (satisfying $v(t,r)=\partial_{r}v_{0}(t,r)$) is identical in both ZSG and uniform-expansion gauge (UEG), the other quantities differ as
$\Phi_{\rm ZSG}=\Phi+\dot{\chi},\qquad\Psi_{\rm ZSG}=\Psi+H\chi,\qquad\delta w_{\rm ZSG}=\delta w-4aHw_{b}\left(v_{0}+\frac{\chi}{a}\right),$ (A.1)
where the quantities in the right-hand side are measured in the UEG. However, we argue that their difference is small in the following sense. On general ground we expect
$\Phi\sim\Psi\sim\frac{\delta w}{w_{b}}\sim v.$ (A.2)
On the other hand, as discussed in Eq. (3.24), we require the condition
$\frac{\partial_{r}\chi}{a}\sim\frac{\chi}{l_{c}}\ll v,$ (A.3)
with $l_{c}$ being the characteristic length scale in the present system. In the regime of our interest, we can identify it with the sound shell thickness. Therefore, the differences in the two gauges behave as
$\dot{\chi}\sim H\chi\ll Hl_{c}v,\qquad aHw_{b}\left(v_{0}+\frac{\chi}{a}\right)\sim aHw_{b}\left(\frac{l_{c}v}{a}+\frac{\chi}{a}\right)\sim Hl_{c}w_{b}v,$ (A.4)
and thus we expect a suppression factor of $Hl_{c}<1$. We finally note that as discussed in Refs. [93, 82], zero-shear gauge results in the inconsistency with the spherical symmetric solution of Tolman-Oppenheimer-Volkoff [94, 95] when the fluid has relativistic pressure. Notice that this issue does not exist in the uniform-expansion gauge, which provides additional motivation for adopting it in our system.
Appendix B True vacuum region as the background
@@
$\delta\tilde{p}$ $=\frac{1}{4}\delta\tilde{w}-\frac{3}{4}\alpha(t)\Theta(r-r_{w}(t)),$ (B.2)
$\delta\tilde{\rho}$ $=\frac{3}{4}\delta\tilde{w}+\frac{3}{4}\alpha(t)\Theta(r-r_{w}(t)).$ (B.3)
@@
Appendix C Late time evolution